A car initially at rest, accelerates at 7.29 m/s to the 2nd power for 11.3 s. How far did it go in this time?


1
Expert's answer
2021-04-09T06:38:48-0400

To be given in question

Car initial velocity "u=0"

"v=7.29 meter\/sec"

"t=11.3sec"

To be asked in question


Distance "s" =?

First law of newton motion

"v=u+at"

Put value

"7.29=0+a\times11.3"

a=0.645"meter\/sec^2"

second law of newton motion

"s=ut+\frac{1}{2}at^2"

"=0+\frac{1}{2}\times0.645\times(11.3)^2"

"s=41.19 meter"